# HESSO: Towards Automatic, Efficient, and User Friendly Any Neural Network Training and Pruning

### v1

@@ 3.1 Saliency Score
After warming up $T_{w}$ steps in Algorithm 1, HESSO has typically collected reasonable information regarding the gradient and the iterate. It then starts to identify redundant groups upon the target group sparsity level $K$ to partition the groups $\mathcal{G}$ into important group set $\mathcal{G}_{I}$ and redundant group set $\mathcal{G}_{R}$, i.e., $\mathcal{G}_{I}\bigcup\mathcal{G}_{R}=\mathcal{G}$ and $|\mathcal{G}_{R}|=K$.
@@
HESSO achieves it by periodically measuring the importance of each parameter group $g\in\mathcal{G}$. To begin, we initialize the important group set as the whole group set $\mathcal{G}_{I}\leftarrow\mathcal{G}$, and the redundant group set as empty $\mathcal{G}_{R}\leftarrow\emptyset$. Given a pre-defined pruning periods $P$, we identify $\widehat{K}\leftarrow K/P$ important groups to designate as redundant during each period. The redundant groups are the ones with bottom-$\widehat{K}$ saliency scores.
$\begin{split}\mathcal{G}_{R}&\leftarrow\mathcal{G}_{R}\bigcup\mathop{{\text{Bottom-}\widehat{K}}}_{g\in\mathcal{G}_{I}}\ \text{SaliencyScore}([\bm{x}]_{g},[\nabla\!f(\bm{x})]_{g})\\ \mathcal{G}_{I}&\leftarrow\mathcal{G}_{I}/\mathop{{\text{Bottom-}\widehat{K}}}_{g\in\mathcal{G}_{I}}\ \text{SaliencyScore}([\bm{x}]_{g},[\nabla\!f(\bm{x})]_{g})\end{split}$ (6)
The selection of the saliency score in HESSO is flexible and can be tailored to different purposes. By default, the following categories are considered:
$\text{score}_{\text{mag}}([\bm{x}]_{g})\leftarrow\left\|[\bm{x}]_{g}\right\|_{2},\ \text{score}_{\text{mag}}([\bm{x}]_{g})\leftarrow\text{score}_{\text{mag}}([\bm{x}]_{g})/\sum_{g\in\mathcal{G}_{I}}\text{score}_{\text{mag}}([\bm{x}]_{g}).$ (7)
$\text{score}_{\text{avg-mag}}([\bm{x}]_{g})\leftarrow\left\|[\bm{x}]_{g}\right\|_{2}/|\sqrt{|g|}|,\ \text{score}_{\text{avg-mag}}([\bm{x}]_{g})\leftarrow\text{score}_{\text{avg-mag}}([\bm{x}]_{g})/\sum_{g\in\mathcal{G}_{I}}\text{score}_{\text{avg-mag}}([\bm{x}]_{g}).$ (8)
$\begin{split}\text{score}_{\text{cosine}}([\bm{x}]_{g},[\nabla\!f(\bm{x})]_{g})&\leftarrow[\bm{x}]_{g}^{\top}[\nabla\!f(\bm{x})]_{g}/(\left\|[\bm{x}]_{g}\right\|\left\|[\nabla\!f(\bm{x})]_{g}\right\|),\\ \text{score}_{\text{cosine}}([\bm{x}]_{g},[\nabla\!f(\bm{x})]_{g})&\leftarrow\text{score}_{\text{cosine}}([\bm{x}]_{g})/\sum_{g\in\mathcal{G}_{I}}\text{score}_{\text{cosine}}([\bm{x}]_{g}).\end{split}$ (9)
Taylor Importance. To further quantitatively approximate the effect of projecting the parameter group $[\bm{x}]_{g}$ onto zero on the objective function, we can employ the Taylor expansion. Taylor expansion could estimate the impact of small changes in the parameters on the function value, allowing us to consider varying orders of Taylor importance. In particular, the first-order Taylor expansion provides a linear approximation of the objective function around the current parameter point. The impact of setting $[\bm{x}]_{g}\to\bm{0}$ can be estimated by the dot product of the gradient and the change in parameters. This approximation helps identify groups whose removal will likely decrease the objective function.
$\begin{split}\text{score}_{\text{Taylor-1st}}([\bm{x}]_{g},[\nabla\!f(\bm{x})]_{g})&\leftarrow|f(\bm{x})-f(\bm{x}|[\bm{x}]_{g}\to\bm{0})|\approx|[\bm{x}]_{g}^{\top}[\nabla\!f(\bm{x})]_{g}|,\\ \text{score}_{\text{Taylor-1st}}([\bm{x}]_{g},[\nabla\!f(\bm{x})]_{g})&\leftarrow\text{score}_{\text{Taylor-1st}}([\bm{x}]_{g},[\nabla\!f(\bm{x})]_{g})/\sum_{g\in\mathcal{G}_{I}}\text{score}_{\text{Taylor-1st}}([\bm{x}]_{g},[\nabla\!f(\bm{x})]_{g}).\end{split}$ (10)
$\begin{split}\text{score}_{\text{Taylor-2nd}}([\bm{x}]_{g},[\nabla\!f(\bm{x})]_{g})&\leftarrow|f(\bm{x})-f(\bm{x}|[\bm{x}]_{g}\to\bm{0})|\approx[\bm{x}]_{g}^{\top}[\nabla\!f(\bm{x})]_{g}+\frac{1}{2}[\bm{x}]_{g}^{\top}[\nabla\!^{2}f(\bm{x})]_{g}[\bm{x}]_{g},\\ \text{score}_{\text{Taylor-2nd}}([\bm{x}]_{g},[\nabla\!f(\bm{x})]_{g})&\leftarrow\text{score}_{\text{Taylor-2nd}}([\bm{x}]_{g},[\nabla\!f(\bm{x})]_{g})/\sum_{g\in\mathcal{G}_{I}}\text{score}_{\text{Taylor-2nd}}([\bm{x}]_{g},[\nabla\!f(\bm{x})]_{g}).\end{split}$ (11)
3.2 Hybrid Training in HESSO
After identifying the redundant groups in Section 3.1, the next step is to project these groups onto zero and transfer their knowledge to the important groups, ensuring the pruned model maintains its performance. This is achieved through a hybrid training schema.
For the redundant groups $\mathcal{G}_{R}$, we progressively and uniformly push their parameters towards zero. This process is detailed in line 3-4 in Algorithm 1 and decipted in Figure 3. The goal is to ensure that the parameters in the redundant groups become zero after $T_{p}$ steps. During this penalization process, there is a risk of forgetting the knowledge contained in the redundant groups, which may manifest as a degradation in the objective function’s value. To mitigate this, we employ a standard optimization method, such as vanilla SGD or its variants like Adam, on the important groups $\mathcal{G}_{I}$. This step aims to continue optimizing the objective function $f$ and preserve the model’s performance despite the pruning of redundant groups. By maintaining the optimization of the important groups, the knowledge lost from the redundant groups can be transferred and compensated for, ensuring that the pruned model remains effective.
Compared to DHSPG in OTOv2 (Chen et al., 2023c), HESSO offers more explicit control over sparsity exploration. The pruning process in HESSO is regulated by the pruning periods $P$ and the period length $T_{P}$, which determine the pace and extent of the pruning procedure. This structured approach simplifies the pruning process, making it easier to manage and understand.
@@ 3.2 Hybrid Training in HESSO
Contrarily, DHSPG involves significant hyper-parameter tuning to adjust parameters related to sparsity exploration. This tuning often requires domain-specific knowledge, as the appropriate settings can vary depending on the particular application or dataset. This requirement can make DHSPG more complex and less accessible, particularly for practitioners without extensive expertise in hyper-parameter and sparse optimization.
@@
As a result, HESSO is generally easier to use and more adaptable to various applications, as it reduces the need for extensive tuning and specialized knowledge. This makes it a more user-friendly option for achieving structured sparsity in models, allowing for more straightforward and consistent application across different tasks and domains.
3.3 Limitations of HESSO
@@
Although HESSO could tackle most DNNs and applications, it may sometimes yield unsatisfactory results when the target DNN possesses certain indispensable structures. The definition of indispensable structure is present as follows.
@@
Given a deep neural network $\mathcal{M}$, a minimally removal structure is called indispensable if removing it from $\mathcal{M}$ would cause performance measurement degrades significantly.
The origin of indispensable structures varies. One reason may be due to architectural design issues where certain layers in $\mathcal{M}$ play more critical roles than others and are very sensitive to any modifications, as exemplified by a low-level vision benchmark in Section 4.2. Another reason could be the learning strategy. For instance, in large language models (LLMs), it has been observed that knowledge is unevenly distributed across different layers (Chen et al., 2023a). Removing any of these structures could result in an irreversible collapse of the DNN’s performance. However, current saliency score identification in HESSO sometimes fails to identify these indispensable components accurately. The underlying reason is that these saliency scores, as discussed in Section 3.1, are designed to approximate the impact of projecting groups of variables to zero. Such approximations, like the commonly used Taylor importance scores, are typically accurate when the iterate is close enough to the origin point. However, during realistic training and pruning, this requirement is usually not met, resulting in unexpected performance issues.
@@ 3.4 Corrective Redundancy Identification Circle
1:Input. Pretraining variable $\bm{x}$, learning rate $\alpha$, termination tolerance $\mathcal{T}$, target group sparsity $K$, sample steps $T$, penalty $\lambda$, and prunable variable partition $\mathcal{G}$.
2:Warm up $\mathcal{B}$ and compute saliency scores.
3:Initialize $\mathcal{S}$ to store saliency scores for each $g\in\mathcal{G}$.
4:Initialize violating group set $\mathcal{V}$
@@
5:Initialize historical set $\mathcal{H}\leftarrow\mathcal{V}$.
6:while $|\mathcal{V}|\leq\mathcal{T}$ do
@@
7:     Initialize trial violating group set $\widehat{\mathcal{V}}\leftarrow\emptyset$.
8:     Initialize $\alpha_{0}\leftarrow\alpha$, $\lambda_{0}\leftarrow\lambda$, and $\bm{x}_{0}\leftarrow\bm{x}$.
9:     for $t=0,1,\cdots,T-1$ do
10:         Compute gradient of $f$ over $\bm{x}_{t}$ as $\nabla\!f(\bm{x}_{t})$.
11:         Compute trial $\tilde{\bm{x}}_{t+1}\leftarrow\bm{x}_{t}-\alpha_{t}\nabla\!f(\bm{x}_{t})$.
12:         Penalize variables in the violating set
$[\bm{x}_{t+1}]_{\mathcal{V}}\leftarrow[\tilde{\bm{x}}_{t+1}]_{\mathcal{V}}-\lambda_{t}[\bm{x}_{t}]_{\mathcal{V}}$ (13)
13:         Compute saliency scores of $\mathcal{G}$ and collect into $\mathcal{S}$.
14:         Update trial set $\widehat{\mathcal{V}}$ if new violating groups appear.
@@
$\widehat{\mathcal{V}}\leftarrow\widehat{\mathcal{V}}\cup\{g:g\in\mathcal{G}\ \text{with bottom-K scores}\}/\mathcal{V}$ (14)
15:         Update penalty $\lambda_{t}$ and learning rate $\alpha_{t}$.
16:     end for
17:     Update violating set $\mathcal{V}\leftarrow\widehat{\mathcal{V}}/\mathcal{H}$.
18:     Update historical set $\mathcal{H}\leftarrow\mathcal{H}\bigcup\mathcal{V}$.
19:end while
20:Set redundant set $\mathcal{G}_{R}$ upon saliency score collection $\mathcal{S}$.
$\mathcal{G}_{R}\leftarrow\{g:g\ \text{with bottom-K scores in}\ \mathcal{S}\}$ (15)
21:Return. Identified redundant group set $\mathcal{G}_{R}$ and important group set $\mathcal{G}_{I}$ as $\mathcal{G}/\mathcal{G}_{R}$.
@@
To mitigate the issue of false positive redundant predictions caused by the approximation error of Taylor expansion, CRIC measures the saliency score of redundant group candidates multiple times along the projection to the origin. Unlike the greedy approach in HESSO, CRIC incorporates a corrective cycle mechanism. This mechanism iteratively promotes groups as redundant and tracks the outlier groups. The cycle terminates when the redundancy prediction is deemed reliable, i.e., no outlier appearance is detected. The final output is a set of redundant groups $\mathcal{G}_{R}$ with the bottom-K overall saliency scores. This approach significantly reduces false positive redundant identifications and addresses the failure cases of HESSO, as demonstrated numerically in Section 4.
@@
In Algorithm 2, we utilize a violating group set $\mathcal{V}$ to track outlier or violating groups, which are more redundant or deviate from the current redundant group prediction. $\mathcal{V}$ is initialized with the group set having the bottom-K saliency scores (see line 4). A historical set $\mathcal{H}$ is also used to track groups whose saliency scores have been fully exploited through multiple sampling along the projection to the origin. This set is initialized as empty $\emptyset$, as shown in line 5.
@@
When the violating set is fairly large, i.e., $|\mathcal{V}|>\mathcal{T}$ with $\mathcal{T}$ as a predefined terminating tolerance, we progressively project these violating groups onto zero. By default, saliency score sampling points are uniformly distributed along the projection process. Groups with lower importance scores that have not been visited in $\mathcal{H}$ are added to a newly constructed violating set $\hat{\mathcal{V}}$ for the next corrective cycle. The corrective cycling algorithm continues until violating instances rarely appear, i.e., $|\mathcal{V}|\leq\mathcal{T}$, see line 6.
@@
Theorem 3.2 guarantees that CRIC terminates within a finite number of iterations, preventing endless loops. Furthermore, Corollary 3.3 provides an upper bound on the number of cycles required by CRIC, ensuring a practical and efficient pruning process.
Algorithm 3 HESSO with CRIC
1:Input. pretraining variable $\bm{x}_{0}$, learning rate $\alpha$.
2:Use CRIC in Algorithm 2 to get redundant and important group sets $\mathcal{G}_{R}$ and $\mathcal{G}_{I}$.
3:Set up regularizer $\lambda$ based on CRIC information.
4:for $t=0,1,\cdots$ do
5:     Compute trial iterate
$\widehat{\bm{x}}_{t+1}\leftarrow\bm{x}_{t}-\alpha_{t}\nabla\!f(\bm{x}_{t})$ (16)
6:     For important groups $\mathcal{G}_{I}$, update the next iterate as the trial iterate.
$[\bm{x}_{t+1}]_{\mathcal{G}_{I}}\leftarrow[\widehat{\bm{x}}_{t+1}]_{\mathcal{G}_{I}}.$ (17)
7:     For each redundant group $g\in\mathcal{G}_{R}$, penalize the trial variable via regularizer
$[\widehat{\bm{x}}_{t+1}]_{g}\leftarrow[\widehat{\bm{x}}_{t+1}]_{g}-\lambda_{g}[\bm{x}_{g}]/\left\|[\bm{x}]_{g}\right\|$ (18)
8:     Perform Half-Space or proximal operator on $\mathcal{G}_{R}$ to yield next iterate over redundant groups.
9:end for
10:Output. The final iterate $\bm{x}^{*}$
Theorem 3.2.
@@
Corollary 3.3.
Given the terminating tolerance $\mathcal{T}$, the CRIC terminates with no more $(|\mathcal{G}|-K)/\max{}\{\mathcal{T},1\}$.
Once the corrective cycles terminate, the saliency scores obtained are deemed reliable. At this point, the redundant set $\mathcal{G}_{R}$ is constructed based on these reliable saliency scores, as indicated in line 20. This set of redundant groups is then returned for further use, such as hybrid training in HESSO, as detailed in Algorithm 2. For simplicity, the HESSO variant that utilizes CRIC for identifying redundant groups is referred to as HESSO-CRIC throughout the paper (outlined in Algorithm 3). This naming convention helps distinguish the variant from the original HESSO method, emphasizing the integration of the corrective cycle mechanism to enhance the reliability of the pruning process.

### v2

@@ 3.1 Saliency Score
After warming up $T_{w}$ steps in Algorithm 1, HESSO has typically collected reasonable information regarding the gradient and the iterate. It then starts to identify redundant groups upon the target group sparsity level $K$ to partition the groups $\mathcal{G}$ into important group set $\mathcal{G}_{I}$ and redundant group set $\mathcal{G}_{R}$, i.e., $\mathcal{G}_{I}\bigcup\mathcal{G}_{R}=\mathcal{G}$ and $|\mathcal{G}_{R}|=K$. HESSO achieves it by periodically measuring the importance of each parameter group $g\in\mathcal{G}$. To begin, we initialize the important group set as the whole group set $\mathcal{G}_{I}\leftarrow\mathcal{G}$, and the redundant group set as empty $\mathcal{G}_{R}\leftarrow\emptyset$. Given a pre-defined pruning periods $P$, we identify $\widehat{K}\leftarrow K/P$ important groups to designate as redundant during each period. The redundant groups are the ones with bottom-$\widehat{K}$ saliency scores. In particular, the redundant group set ${\cal G}_{R}$ and the important group set ${\cal G}_{I}$ are updated as follows:
$\begin{split}&\mathcal{G}_{R}\leftarrow\mathcal{G}_{R}\bigcup\mathop{{\text{Bottom-}\widehat{K}}}_{g\in\mathcal{G}_{I}}\ \text{SaliencyScore}([\bm{x}]_{g},[\nabla\!f(\bm{x})]_{g}),\\ &\mathcal{G}_{I}\leftarrow\mathcal{G}_{I}/\mathop{{\text{Bottom-}\widehat{K}}}_{g\in\mathcal{G}_{I}}\ \text{SaliencyScore}([\bm{x}]_{g},[\nabla\!f(\bm{x})]_{g}).\end{split}$
The selection of the saliency score in HESSO is flexible and can be tailored to different purposes. By default, we consider the categories presented in Appendix B.
3.2 Hybrid Training in HESSO
After identifying the redundant groups in Section 3.1, the next step involves projecting these groups onto zero and transfering their knowledge to the important groups, ensuring that the pruned model retains its performance. This is accomplished through a hybrid training schema.
For the redundant groups $\mathcal{G}_{R}$, we progressively and uniformly push their parameters towards zero. This process is detailed in line 11-12 in Algorithm 1 and decipted in Figure 3. The goal is to ensure that the parameters in the redundant groups become zero after $T_{p}$ steps. During this penalization process, there is a risk of forgetting the knowledge contained in the redundant groups, which may manifest as a degradation in the objective function’s value. To mitigate this, we employ a standard optimization method, such as vanilla SGD or its variants such as Adam, on the important groups $\mathcal{G}_{I}$. This step aims to continue optimizing the objective function $f$ and preserve the model’s performance despite the pruning of redundant groups. By maintaining the optimization of the important groups, the knowledge lost from the redundant groups can be transferred and compensated for, ensuring that the pruned model remains effective.
@@ 3.2 Hybrid Training in HESSO
Next, we provide brief intuitive comparisons of HESSO against two popular pruning algorithms.
Minimize tuning efforts compared to DHSPG.
@@
DHSPG in OTOv2 involves significant hyper-parameter tuning to adjust parameters for sparsity exploration. This tuning often requires domain-specific knowledge, as the appropriate settings can vary depending on the particular application or dataset. This requirement can make DHSPG more complex and less accessible, particularly for practitioners without extensive expertise in hyper-parameter and sparse optimization. Contrarily, HESSO offers more explicit control over sparsity exploration. The pruning process in HESSO is regulated by the pruning periods $P$ and the period length $T_{P}$, which determine the pace and extent of the pruning procedure. This structured approach simplifies the process, making it easier to manage.
Architecture-agnostic computational invariance compared to ResRep and SliceGPT.
ResRep (Ding et al., 2021b) and SliceGPT (Ashkboos et al., 2024) are proposed to preserve computational invariance, i.e., making pruned and full models produce similar outputs, for CNNs and transformers, respectively. However, they are architecture specific, requires additional efforts, such as injecting additional layers in SliceGPT and computing reset gradients in ResRep. The knowledge transfer in HESSO similarly seeks to maintain computational invariance but does so by preserving objective function levels. In addition, HESSO is architecture-agnostic, efficient and user-friendly, demonstrating both scalability and versatility compared with ResRep and SliceGPT.
@@
As a result, HESSO is generally easier to use and more adaptable to various applications, as it significantly reduces the need for extensive tuning and specialized knowledge. The design of hybrid training for knowledge transfer effectively promotes the performance of pruned model. It makes HESSO a more efficient and user-friendly option for achieving structured sparsity and ensuring consistent application across different tasks and domains.
3.3 Approximation Errors of Saliency Scores
@@
Although HESSO can tackle most DNNs and tasks, it may yield unsatisfactory results when the target DNN possesses certain indispensable structures, defined as follows.
@@
Given a deep neural network $\mathcal{M}$, a minimally removal structure is called indispensable if removing it from $\mathcal{M}$ would cause significant performance degradation, which can not be recovered given user resources. In particular, we say a minimally removal structure as $\epsilon$-indispensable associated with an objective $f$ if pruning the variables $[\bm{x}]_{g}\to\bm{0}$ deteriorates $f$ at least $\epsilon$, i.e., $f(\bm{x}|[\bm{x}]_{g}\to\bm{0})\geq f(\bm{x})+\epsilon$ for a minimization optimization problem. The degradation $\epsilon$ can not be recovered by (i) keeping training $\mathcal{M}$, (ii) the training cost such as GPU days exceeding user budget, or (iii) the training receipt for $\mathcal{M}$ is black-box and hard to be reproduced.
The origin of indispensable structures varies. One reason may be due to architectural design issues where certain layers in $\mathcal{M}$ play more critical roles than others and are very sensitive to any modifications, as exemplified by a low-level vision benchmark in Section 4.2. Another reason could be the learning strategy. For instance, in large language models (LLMs), it has been observed that knowledge is unevenly distributed across different layers (Chen et al., 2023a). Removing any of these structures could result in an irreversible collapse of the DNN’s performance.
Saliency score approximation errors.
The existing saliency scores might fail to identify these indispensable components accurately. As described in Appendix B, they are typically designed to approximate the impact of projecting groups of variables to zero over the objective function. Such approximations, for example, perhaps the most commonly used Taylor importance scores, are more accurate when the iterate is close enough to the origin point.
Theorem 3.2 (Approximation error of Taylor importance).
Suppose the gradient and second-order derivative of $f$ are bounded. Use first-order $m^{L}$ and second-order $m^{Q}$ Taylor approximations to measure the function value $f$ after pruning $g\in\mathcal{G}$, i.e., $[\bm{x}]_{g}\to\bm{0}$. Let $\bm{s}$ satisfy $[\bm{s}]_{\mathcal{G}/g}=[\bm{0}]_{\mathcal{G}/g}$ and $[\bm{s}]_{g}=-[\bm{x}]_{g}$, Then the approximation error bound $|f(\bm{x}+\bm{s})-m^{L}(\bm{x}+\bm{s})|$ and $|f(\bm{x}+\bm{s})-m^{Q}(\bm{x}+\bm{s})|$ are proportional to $\mathcal{O}(\left\|[\bm{x}]_{g}\right\|^{2})$ and $\mathcal{O}(\left\|[\bm{x}]_{g}\right\|^{3})$, respectively.
However, during realistic training and pruning, this requirement is usually not met. As stated in Theorem 3.2, the approximation error bounds increase proportionally with $\left\|[\bm{x}]_{g}\right\|$, indicating that the further the distance from the origin, the larger the approximation error. As a result, this can lead to the false positively pruning of indispensable structures, which in turn causes performance issues.
@@ 3.4 Corrective Redundancy Identification Circle
1:Input. Trainable variable $\bm{x}$, learning rate $\alpha$, termination tolerance $\mathcal{T}$, target group sparsity $K$, sample steps $T$, and prunable variable partition $\mathcal{G}$.
2:Initialize $\mathcal{S}$ to store saliency scores for each $g\in\mathcal{G}$.
3:Initialize violating group set $\mathcal{V}$:
@@
4:Initialize historical set $\mathcal{H}\leftarrow\mathcal{V}$.
5:while $|\mathcal{V}|\leq\mathcal{T}$ do
@@
6:     Initialize trial violating group set $\widehat{\mathcal{V}}\leftarrow\emptyset$.
7:     Initialize $\alpha_{0}\leftarrow\alpha$, $\lambda_{0}\leftarrow\lambda$, and $\bm{x}_{0}\leftarrow\bm{x}$.
8:     for $t=0,1,\cdots,T-1$ do
9:         Compute trial $\tilde{\bm{x}}_{t+1}\leftarrow\bm{x}_{t}-\alpha_{t}\nabla\!f(\bm{x}_{t})$.
10:         Penalize variables in the violating set:
$[\bm{x}_{t+1}]_{\mathcal{V}}\leftarrow\frac{T-t-1}{T-t}\frac{[\bm{x}_{t}]_{\mathcal{V}}}{{\left\|[\tilde{\bm{x}}_{t+1}]_{\mathcal{V}}\right\|}}.$
11:         Compute saliency scores of $\mathcal{G}$ and merge to $\mathcal{S}$.
12:         Update set $\widehat{\mathcal{V}}$ if new violating groups appear:
@@
$\quad\quad\quad\widehat{\mathcal{V}}\leftarrow\widehat{\mathcal{V}}\cup\{g:g\in\mathcal{G}\ \text{with bottom-K scores}\}/\mathcal{V}.$
13:         Update penalty $\lambda_{t}$ and learning rate $\alpha_{t}$.
14:     end for
15:     Update violating set $\mathcal{V}\leftarrow\widehat{\mathcal{V}}/\mathcal{H}$.
16:     Update historical set $\mathcal{H}\leftarrow\mathcal{H}\bigcup\mathcal{V}$.
17:end while
18:Set redundant set $\mathcal{G}_{R}$ upon saliency score collection $\mathcal{S}$:
$\mathcal{G}_{R}\leftarrow\{g:g\ \text{with bottom-K scores in}\ \mathcal{S}\}.$
19:Return. Identified redundant group set $\mathcal{G}_{R}$ and important group set $\mathcal{G}_{I}$ as $\mathcal{G}/\mathcal{G}_{R}$.
@@
To mitigate the issue of false positive redundant predictions caused by the approximation error, such as Taylor expansion, CRIC measures the saliency score of redundant group candidates multiple times along the projection to the origin. Unlike the greedy approach in HESSO, CRIC incorporates a corrective cycle mechanism. This mechanism iteratively promotes groups as redundant and tracks the outlier groups. The cycle terminates when the redundancy prediction is deemed reliable, i.e., no outlier appearance is detected. The final output is a set of redundant groups $\mathcal{G}_{R}$ with the bottom-K overall saliency scores. This approach significantly reduces false positive redundant identifications and addresses the failure cases of HESSO, as demonstrated numerically in Section 4.
@@
In Algorithm 2, we utilize a violating group set $\mathcal{V}$ to track outlier or violating groups, which are more redundant or deviate from the current redundant group prediction. $\mathcal{V}$ is initialized with the group set having the bottom-K saliency scores (see line 3). A historical set $\mathcal{H}$ is also used to track groups whose saliency scores have been fully exploited through multiple sampling along the projection to the origin. This set is initialized as empty $\emptyset$, as shown in line 4.
@@
When the violating set is fairly large, i.e., $|\mathcal{V}|>\mathcal{T}$ with $\mathcal{T}$ as a predefined terminating tolerance which is by default as empty set, i.e., $\mathcal{T}=\emptyset$, we progressively project these violating groups onto zero. By default, saliency score sampling points are uniformly distributed along the projection process. Groups with lower importance scores that have not been visited in $\mathcal{H}$ are added to a newly constructed violating set $\hat{\mathcal{V}}$ for the next corrective cycle. The corrective cycling algorithm continues until violating instances rarely appear, i.e., $|\mathcal{V}|\leq\mathcal{T}$, see line 5.
@@
Theorem 3.3 guarantees that CRIC terminates within a finite number of iterations, preventing endless loops and executing efficiently. We provided detailed proof for Theorem 3.3 in Appendix A. Furthermore, Corollary 3.4 provides an upper bound on the number of cycles required by CRIC, ensuring a practical and efficient pruning process.
Theorem 3.3 (Finite termination of CRIC).
@@
Corollary 3.4 (Upper bounds of cycle numbers).
Given the terminating tolerance $\mathcal{T}$, the CRIC terminates with no more $(|\mathcal{G}|-K)/\max{}\{\mathcal{T},1\}$ cycles.
Once the corrective cycles terminate, the saliency scores obtained are deemed reliable. At this point, the redundant set $\mathcal{G}_{R}$ is constructed based on these reliable saliency scores, as indicated in line 18. This set of redundant groups is then returned for further use, such as hybrid training in HESSO (as detailed in Algorithm 1). For simplicity, the HESSO variant that utilizes CRIC for identifying redundant groups is referred to as HESSO-CRIC throughout the paper, as outlined in Algorithm 3. This naming convention distinguishes the variant from HESSO, highlighting the addition of the corrective cycle mechanism that enhances the reliability of the pruning process.
Algorithm 3 HESSO-CRIC
1:Input. trainable variable $\bm{x}_{0}$, learning rate $\alpha$, warm-up steps, $T_{w}$, and hybrid training steps $T_{h}$.
2:Warm-up for $T_{w}$ steps via SGD or its variants.
3:Use CRIC in Algorithm 2 to get redundant and important group sets $\mathcal{G}_{R}$ and $\mathcal{G}_{I}$.
4:Hybrid Training for Knowledge Transfer.
5:for $t=0,1,\cdots,T_{h}$ do
6:     Compute trial iterate $\widehat{\bm{x}}_{t+1}\leftarrow\bm{x}_{t}-\alpha_{t}\nabla\!f(\bm{x}_{t})$.
7:     Compute transferring penalty ratio $[\bm{\gamma}_{t}]_{g}\leftarrow\frac{T-t-1}{T-t}\frac{\left\|[\bm{x}_{t}]_{g}\right\|}{\left\|[\widehat{\bm{x}}_{t+1}]_{g}\right\|}$ for each $g\in{\mathcal{G}}_{R}$.
8:     Update redundant group variables $[\bm{x}_{t+1}]_{{\mathcal{G}}_{R}}\leftarrow[\bm{\gamma}_{t}]_{{\mathcal{G}}_{R}}[\widehat{\bm{x}}_{t+1}]_{{\mathcal{G}}_{R}}.$
9:     Update important group variables $[\bm{x}_{t+1}]_{\mathcal{G}_{I}}\leftarrow[\widehat{\bm{x}}_{t+1}]_{\mathcal{G}_{I}}$.
10:end for
11:Keep training variables in important groups till convergence.
12:Output. The final iterate $\bm{x}^{*}$.